# Frequentist Validity of Epistemic Uncertainty Estimators

### v1

@@ 3.4 Distinguishing Algorithmic Randomness from Statistical Uncertainty
Deep neural networks are generally trained with variants of stochastic gradient descent and a random initialization. As such, the fitted model is random even with the training data fixed. In this section, we show that our bootstrap estimator provides a way to quantify how much of the total variation in the fitted model is due to randomness in the algorithm versus statistical fluctuations in the training set.
@@
$=$ $\mathrm{H}(\mathbb{E}_{(\mathcal{D}_{n}^{b},s)|\mathcal{D}_{n}}[\hat{p}(\mathcal{D}_{n}^{b},s)])-\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])]+\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])]-\mathbb{E}_{(\mathcal{D}_{n}^{b},s)|\mathcal{D}_{n}}[\mathrm{H}(\hat{p}(\mathcal{D}_{n}^{b},s))]$
$=$ $\underbrace{\mathrm{H}(\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)]])-\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])]}_{=:\mathrm{I}^{\textrm{resampling}}_{b}}+\underbrace{\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}\left[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])-\mathbb{E}_{s}[\mathrm{H}(\hat{p}(\mathcal{D}_{n}^{b},s))]\right]}_{=:\mathrm{I}^{\textrm{seeds}}_{b}}.$ (3)
$\mathrm{I}^{\textrm{resampling}}_{b}$ can be viewed as an estimator of the epistemic uncertainty arising due to the randomness in data drawing. On the other hand, $\mathrm{I}^{\textrm{seeds}}_{b}$ captures the amount of the entropy difference explained by the variation in the random seeds. That is, this term represents the amount of variation due to the algorithmic randomness. Thus, our proposed estimator can be decomposed into a ‘data sampling’ component and a ‘training randomness’ component.
It is particularly interesting to connect this decomposition with deep ensembles. Notice that the estimate of the mutual information from deep ensembles is
$\mathrm{I}^{\textrm{deep ensemble}}:=$ $\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n},s)])-\mathbb{E}_{s}[\mathrm{H}(\hat{p}(\mathcal{D}_{n},s))].$
This is parallel to the term $\mathrm{I}^{\textrm{seeds}}_{b}$ above, but with the full data instead of a bootstrap subsample. Thus, the bootstrap estimator above can be viewed as approximately the MI from random seeds plus an additional term capturing variability in the dataset. In the experiments, we will see that the latter is substantially smaller than the former. In this way, $\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})\approx\mathrm{I}^{\textrm{deep ensemble}}$. This link gives a new frequentist motivation for deep ensembles.
@@ 4 Experiments
4.1 Validation of the Plug-In Principle
We begin our empirical investigations with validation of the “plug-in” principle—specifically whether the variance under true redraws of the data matches the variance over the bootstrapped datasets.
Since our proposed estimator relies on $\hat{p}(X_{\mathrm{test}};\theta)$, we choose to directly investigate how much variation occurs in $\hat{p}(X_{\mathrm{test}};\theta)$ when we train on multiple datasets drawn from the same population.
Experimental Setup: We use the “The Street View House Numbers” (SVHN) Dataset (Netzer et al.,, 2011) which consist of images of digits of house-numbers, along with the ResNet-18 (He et al.,, 2015) architecture. We defer hyperparameter details to Appendix B.1. We first select at random 50,000 training images from the SVHN dataset (denote this by $\mathcal{D}_{0}$) and train a deep ensemble of 10 models. Let us denote these models by $f^{(1)}_{0},...,f^{(10)}_{0}$ and the deep ensemble model by $\bar{f}_{0}$ (the predictions of the ensemble model are simply the average of constituent model predictions). Next, we select 10 more training datasets each of 50,000 images from the SVHN dataset (denote these by $\mathcal{D}_{1},...,\mathcal{D}_{10}$) such that all the training datasets are disjoint and represent i.i.d. draws of the training data from the same population. We then train a deep ensemble on each of these datasets too (denote the ensemble models by $\bar{f}_{1},...,\bar{f}_{10}$).
Figure 2: (a) Spread under Redraws of Data: Scatter plot of standard deviation in the probability prediction corresponding to the true class for models trained on redraws of the training dataset versus (left) models trained on bootstrapped versions of one dataset and (right) models within a deep ensemble along with Pearson ($\rho$) and Spearman ($r_{s}$) coefficients. See Sec. 4.1 for details. (b) Active Learning: Test accuracy on MNIST for four data acquisition functions. We plot the mean and shade 2 standard deviations across 10 repetitions.
Investigations: To estimate the spread in predictions from $\bar{f}_{0}$, we calculate the standard deviation in the probability predictions corresponding to the true class obtained from $\bar{f}_{1}$ to $\bar{f}_{10}$ for each test instance in the SVHN dataset.
To compare this to the bootstrap estimate, we create 10 bootstrapped datasets from $\mathcal{D}_{0}$ and train a deep ensemble on each of these (denote these by $\bar{f}_{b1},...,\bar{f}_{b10}$). We similarly calculate the standard deviation in the probability predictions corresponding to the true class obtained from $\bar{f}_{b1}$ to $\bar{f}_{b10}$ for each test instance.
Fig. 2(a)(left) shows a scatter plot of the standard deviation under redraws and under bootstrapping for each test instance. As we can see, the two correlate well with a Pearson coefficient of $\rho=0.84$ and a Spearman rank correlation coefficient of $r_{s}=0.86$. The standard deviations under redraws are slightly higher as may be expected due to more variability in training data.
We also use this experimental setup to begin our study into the performance of deep ensembles from a frequentist lens. Specifically, we investigate the spread within a deep ensemble by calculating the standard deviation in the probability prediction corresponding to the true class obtained from $f^{(1)}_{0}$ to $f^{(10)}_{0}$ for each test instance and visualize this against the standard deviation across $\bar{f}_{1}$ to $\bar{f}_{10}$ in Fig. 2(a)(right).
The spread within a deep ensemble does not correlate as well with the spread across redraws of the data as the bootstrap spread does. The spread within a deep ensemble is often significantly higher than spread under redraws of the data. This illustrates an important but often ignored point—variability in ensembled predictions is not the same as variability within the ensemble. If practitioners are interested in estimating the frequentist variability of their ensemble prediction, it may be better to use bootstrap rather than looking at variability within the ensemble.
4.2 Active Learning
Perhaps the most celebrated use of epistemic uncertainty is active learning and accordingly we study the performance of our measure for such a task. Specifically, we consider a problem setup wherein an algorithm must select points to label from a pool of unlabeled points in order to train a model with high predictive accuracy.
Experimental Details: We follow the exact same experimental procedure as in Smith et al., (2023). Specifically, we use the MNIST dataset (LeCun et al.,, 1998) and train a two layer convolutional neural network with Dropout layers. We train the model for up to 50000 steps of full batch gradient descent with early stopping based on validation set performance. We provide exact hyperparameter details in Appendix B.2.
We initialize the labeled data with just 2 samples from each of the 10 classes and form a pool dataset with 4000 samples from each class from which the algorithm can acquire labels. Active learning proceeds by training the model on the currently labeled dataset, estimating epistemic uncertainty over all the instances in the pool dataset and acquiring the label for the one with the highest epistemic uncertainty. We do this for up to 300 data acquisitions and track the test accuracy.
We use three different ways of estimating epistemic uncertainty. The most commonly used approach is to leverage the interpretation of Monte Carlo Dropout as Bayesian inference (Gal and Ghahramani,, 2016) wherein Dropout is enabled at inference time and stochastic forward passes are performed to obtain samples from the Bayesian predictive distribution. We obtain 100 such samples for our experiment. Next, we use our frequentist estimate of epistemic uncertainty by running Algorithm 1 for $B=5$. Continuing our investigation into the uncertainty quantification abilities of deep ensembles, we also use the spread within the ensemble to guide data acquisition. Specifically, we use $\mathrm{I}^{\textrm{deep ensemble}}$ (Sec. 3.4) which we estimate by using the set of model weights in the deep ensemble as $\Theta_{b}$ in Algorithm 1. We use a deep ensemble of size 5. Finally, as a baseline we consider completely random acquisition of labels.
Results: Fig. 2(b) plots the mean and shades two standard deviations across 10 replications of the experiment. All three methods of epistemic uncertainty estimation perform similarly, and much better than random acquisition. The success of Dropout in generating posterior samples for mutual information estimation for active learning is well studied and justified through its Bayesian interpretation. The performance of our bootstrap based measure being comparable to Dropout shows its practical success beyond the theoretical justification from Theorem 2. The success of deep ensembles may be expected from their demonstrated capabilities on other uncertainty quantification tasks, but a theoretical justification is lacking.
4.3 Connection with Deep Ensembles
This section is aimed at investigating the success of deep ensembles at uncertainty quantification. In particular, we show in an experiment that the bootstrap MI estimate is approximately equal to that from a deep ensemble. This happens when the dominant contribution to the variation is from the random seeds rather than from statistical variability in the sample, as formalized in Section 3.4.
Experimental Setup: We use a ResNet-18 architecture and the CIFAR-10 dataset (Krizhevsky,, 2009). We defer additional hyperparameter details to Appendix B.3. We first train a deep ensemble consisting of 10 models on the entire dataset. Then, we create 10 bootstrapped datasets and train a deep ensemble of 10 models on each. Recall the notation from Section 3.4. This set of $10\times 10$ bootstrap models allows us to compute $\mathrm{I}^{\textrm{resampling}}_{b}$ and $\mathrm{I}^{\textrm{seeds}}_{b}$ by using empirical averages to approximate the expectations in Eq. 3. We similarly calculate $\mathrm{I}^{\textrm{deep ensemble}}$ as in Sec. 4.2 using the complete data deep ensemble.
Results: Fig. 3(a)(left) shows a scatter plot of $\mathrm{I}^{\textrm{seeds}}_{b}$ versus $\mathrm{I}^{\textrm{resampling}}_{b}$. The dominant contribution to the total mutual information comes from $\mathrm{I}^{\textrm{seeds}}_{b}$ with the line of best fit (red) revealing that it is 5 times $\mathrm{I}^{\textrm{resampling}}_{b}$ on average. Fig. 3(a)(right) shows that $\mathrm{I}^{\textrm{seeds}}_{b}$ and $\mathrm{I}^{\textrm{deep ensemble}}$ correlate quite well with $\rho=0.87,r_{s}=0.90$. This is our key insight—the epistemic uncertainty calculated by considering models within a deep ensemble correctly targets the dominant contribution to our (asymptotically equivalent) estimate of the Bayesian epistemic uncertainty.
Figure 3: (a) Epistemic Uncertainty Decomposition: Scatter plot of epistemic uncertainty due to training stochasticity in the bootstrap procedure ($\mathrm{I}^{\textrm{seeds}}_{b}$) versus (left) epistemic uncertainty due to data resampling in the bootstrap procedure ($\mathrm{I}^{\textrm{resampling}}_{b}$) and (right) deep ensemble estimate of epistemic uncertainty ($\mathrm{I}^{\textrm{deep ensemble}}_{b}$). See Sec. 4.3 for details. (b) Influence Function Approximation: Estimates of epistemic uncertainty using Dropout, versus using influence functions to approximate predictions from bootstrapped datasets.
4.4 Data Attribution Methods
Next, we connect our work to data-attribution methods such as influence functions (Hampel et al.,, 2011; Koh and Liang,, 2017; Giordano et al.,, 2019) and datamodels (Ilyas et al.,, 2022; Park et al.,, 2023) which aim to identify training data points most responsible for a model’s predictions. These methods allow for estimation of predictions under perturbation of training data and can therefore be used to approximate the predictions of a model trained on bootstrapped data needed for our proposed measure. Any data attribution algorithm that yields a mapping from training data to model predictions can be leveraged to estimate the variance terms in (2), and hence give epistemic uncertainty estimates. In this section, we investigate the performance of such an MI estimate based on influence functions. See Appendix B.4 for a similar experiment with datamodels.
Experimental Setup: We train a single convolutional network called the LeNet-5 (LeCun et al.,, 1998) on MNIST data. Next, we calculate the influence functions for each data point as described in Koh and Liang, (2017) for the model parameters in the last two fully connected layers. An influence function for a training data point is an estimate of how much the model parameters change under an infinitesimally small change in the weight of the training point. For each data point, we multiply the influence function by the change in the weighting under bootstrap and add all these products to the parameters to obtain an approximation for model parameters under a bootstrap reweighting. We obtain predictions from 100 such approximated models and use Algorithm 1 to obtain MI estimates. We then use Monte Carlo Dropout to generate posterior 100 predictions as described in Sec. 4.2 for the calculation of MI.
Results: Fig. 3(b) shows the influence function approximation and Dropout estimates. While there is correlation, influence function estimates are an order of magnitude smaller. We hypothesize that this is because the influence function approximation only uses a single trained model and hence fails to capture the randomness due to training stochasticity. Nonetheless, this experiment serves as a proof-of-concept for the potential of data attribution methods in uncertainty quantification.

### v2

@@ 3.4 Distinguishing Algorithmic Randomness from Statistical Uncertainty
Deep neural networks are generally trained with variants of stochastic gradient descent and a random parameter initialization. As such, the fitted model is random even with the training data fixed. Deep ensembles exploit this phenomenon, effectively capturing only the epistemic uncertainty arising from training stochasticity. The bootstrap estimator in Sec. 3.3 provides the crucial conceptual link between the ad-hoc operational technique of deep ensembles and principled Bayesian inference. While standard Bayesian formulations cannot account for stochasticity in model training, in this section we show how the bootstrap estimator naturally decomposes into distinct terms for data variability and training stochasticity, providing a mechanism to isolate the latter. This decomposition will then allow us to study deep ensembles by directly comparing them to the training stochasticity component of a theoretically grounded estimator.
@@
$=\mathrm{H}(\mathbb{E}_{(\mathcal{D}_{n}^{b},s)|\mathcal{D}_{n}}[\hat{p}(\mathcal{D}_{n}^{b},s)])-\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])]$
$+\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])]-\mathbb{E}_{(\mathcal{D}_{n}^{b},s)|\mathcal{D}_{n}}[\mathrm{H}(\hat{p}(\mathcal{D}_{n}^{b},s))]$
$=\underbrace{\mathrm{H}(\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)]])-\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])]}_{=:\mathrm{I}^{\textrm{resampling}}_{b}}$
$+\underbrace{\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}\left[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])-\mathbb{E}_{s}[\mathrm{H}(\hat{p}(\mathcal{D}_{n}^{b},s))]\right]}_{=:\mathrm{I}^{\textrm{seeds}}_{b}}.$ (3)
$\mathrm{I}^{\textrm{resampling}}_{b}$ can be viewed as an estimator of the epistemic uncertainty arising due to the randomness in the data sampling process. Intuitively, this represents how much the model prediction would change if different training data had been sampled from the true distribution. On the other hand, $\mathrm{I}^{\textrm{seeds}}_{b}$ captures the amount of the entropy difference explained by the variation in the random seeds. That is, this term represents the amount of variation due to the algorithmic randomness. Thus, our proposed estimator can be decomposed into a ‘data sampling’ component and a ‘training randomness’ component.
Next, notice that the estimate of the mutual information from deep ensembles is
$\mathrm{I}^{\textrm{deep ensemble}}:=$ $\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n},s)])-\mathbb{E}_{s}[\mathrm{H}(\hat{p}(\mathcal{D}_{n},s))].$ (4)
This is parallel to the term $\mathrm{I}^{\textrm{seeds}}_{b}$ above, but with the full data instead of a bootstrap subsample. Thus, the bootstrap estimator above can be viewed as approximately the MI from random seeds plus an additional term capturing variability in the dataset. In the experiments, we will see that the latter is often smaller than the former. In this way,
$\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})\approx\mathrm{I}^{\textrm{deep ensemble}}.$
This link gives a new frequentist motivation for deep ensembles. More precisely, since Sec. 3.3 proved that the bootstrap estimator is asymptotically correct for the true Bayesian target, the above relation indicates how deep ensembles succeed in capturing the Bayesian MI.
@@ 4 Experiments
Figure 2: Uncertainty Components of Bootstrap Estimator: Scatter plot of epistemic uncertainty due to training stochasticity ($\mathrm{I}^{\textrm{seeds}}_{b}$) versus epistemic uncertainty due to data resampling ($\mathrm{I}^{\textrm{resampling}}_{b}$) along with line of best fit in red. See Sec. 4.2 for details.
Figure 3: Uncertainty due to Training Stochasticity: Scatter plot of epistemic uncertainty due to training stochasticity ($\mathrm{I}^{\textrm{seeds}}_{b}$) versus deep ensemble estimate of epistemic uncertainty ($\mathrm{I}^{\textrm{deep ensemble}}_{b}$) along with line of best fit (orange) and spearman correlation ($\rho$). See Sec. 4.2 for details.
Section 3.4 gives us a way to measure contributions to epistemic uncertainty from sampling variability in the training data and from stochasticity in the training procedure; as well as a measure for the epistemic uncertainty in a deep ensemble. Our empirical investigations now seek to address two questions—
  1. 1.
    Does training randomness ($\mathrm{I}^{\textrm{seeds}}_{b}$ in eq. 3) contribute more to epistemic uncertainty than data sampling ($\mathrm{I}^{\textrm{resampling}}_{b}$ in eq. 3)?
  2. 2.
    Does the deep ensemble ($\mathrm{I}^{\textrm{deep ensemble}}$ in eq. 4) capture the training randomness ($\mathrm{I}^{\textrm{seeds}}_{b}$)?
4.1 Experimental Setup
We will use the same experimental setup across various deep learning classification tasks. We first train a deep ensemble of $B$ models on the entire training dataset. We use this to estimate $\mathrm{I}^{\textrm{deep ensemble}}$ for each test point by approximating expectations with empirical averages in eq. 4. Next, we construct $B$ bootstrapped versions of the training dataset and train a deep ensemble of size $B$ on each. We use these deep ensembles to similarly estimate $\mathrm{I}^{\textrm{seeds}}_{b}$ and $\mathrm{I}^{\textrm{resampling}}_{b}$ for each test point.
4.2 Evaluation Metrics
To answer the first question, we look at the slope of the line of best fit ($m$ when $y=mx$) that regresses $\mathrm{I}^{\textrm{resampling}}_{b}$ on $\mathrm{I}^{\textrm{seeds}}_{b}$; we shall refer to this slope as $m_{\textrm{seeds}}^{\textrm{resampling}}$. Fig. 2 visualizes this line in red. This tells us how the magnitude of $\mathrm{I}^{\textrm{resampling}}_{b}$ compares to $\mathrm{I}^{\textrm{seeds}}_{b}$ on average. For the second question, we again look at the slope of the line of best fit that regresses $\mathrm{I}^{\textrm{deep ensemble}}$ on $\mathrm{I}^{\textrm{seeds}}_{b}$ to compare the relative magnitudes of the two; we shall refer to this slope as $m_{\textrm{seeds}}^{\textrm{deep ensemble}}$. Fig. 3 (right) visualizes this line in orange. We also look at the Spearman rank correlation between $\mathrm{I}^{\textrm{deep ensemble}}$ and $\mathrm{I}^{\textrm{seeds}}_{b}$ to evaluate if points with high relative epistemic uncertainty due to training stochasticity also exhibit high relative epistemic uncertainty when we use deep ensembles; we shall refer to this correlation as $\rho$. This is particularly insightful since performance on many downstream tasks (such as AUC-ROC in outlier detection tasks) often only depends on the ordering of test points rather the explicit magnitudes of the uncertainty estimates.
4.3 Deep Learning Tasks
To lend robustness to our empirical findings, we study performance across four tasks spanning various data domains (natural images, medical images, text), different dataset sizes ($1000$ classes and >1>1M training examples, $2$ classes and 1000 training examples) and different deep learning regimes (training from scratch, fine-tuning, transfer learning). For each task, we defer hyperparameter details and accuracy metrics to Appendix B.1.
4.3.1 CIFAR-10
We train a ResNet-18 (He et al., 2015) from scratch on the CIFAR-10 dataset (Krizhevsky, 2009) which consists of 50,000 training examples spanning 10 classes of natural images. We use ensembles of size 10.
4.3.2 Imagenet
We train a ResNet-18 (He et al., 2015) from scratch on the ImageNet dataset (Deng et al., 2009) which consists of 1.28M training examples spanning 1000 classes of natural images. Due to computational limitations, we use ensembles of size 4.
4.3.3 ISIC Skin Lesion Dataset
The ISIC (International Skin Imaging Collaboration) 2016 skin lesion challenge (Gutman et al., 2016) consists of a dataset of images of skin lesions. There are only 900 training images spanning two classes (benign and malignant) in this dataset. For such small datasets, transfer learning (Pan and Yang, 2010) is a common approach wherein a pretrained model from a similar task is used to extract features for the task of interest. More precisely, we use the pretrained ResNet-50 model from (He et al., 2015) to extract 2048-dimensional feature vectors. We then train a two layer MLP from random initialization on top of these feature vectors to perform binary classification. We use ensembles of size 10.
4.3.4 AG News Dataset
The AG news topic classification dataset (Zhang et al., 2015) consists of 120,000 news articles (titles and descriptions) spanning 4 categories. As is common in natural language processing (NLP), we finetune a pretrained model to perform classification. Specifically, we finetune the TinyBERT (Jiao et al., 2020) model which has a transformer architecture (14.5M parameters) for classification by adding a linear layer (initialized randomly) on top of the transformer layers to perform classification. We use ensembles of size 10.
4.4 Results
Table 1: Slope of line of best fit from $\mathrm{I}^{\textrm{resampling}}_{b}$ on $\mathrm{I}^{\textrm{seeds}}_{b}$ ($m_{\textrm{seeds}}^{\textrm{resampling}}$), from $\mathrm{I}^{\textrm{deep ensemble}}_{b}$ on $\mathrm{I}^{\textrm{seeds}}_{b}$ ($m_{\textrm{seeds}}^{\textrm{deep ensemble}}$) and spearman correlation between $\mathrm{I}^{\textrm{deep ensemble}}_{b}$ and $\mathrm{I}^{\textrm{seeds}}_{b}$ ($\rho$). See Sec. 4.2 for details.
Dataset $m_{\textrm{seeds}}^{\textrm{resampling}}$ $m_{\textrm{seeds}}^{\textrm{deep ensemble}}$ $\rho$
CIFAR-10 0.21 0.94 0.90
ImageNet 0.50 0.91 0.87
ISIC 0.55 0.83 0.91
AG News 0.42 0.88 0.96
Fig. 2 visualizes scatter plots of $\mathrm{I}^{\textrm{seeds}}_{b}$ versus $\mathrm{I}^{\textrm{resampling}}_{b}$ along with the line of best fit (slope = $m_{\textrm{seeds}}^{\textrm{resampling}}$) with numerical values of the slopes in Table 1. $\mathrm{I}^{\textrm{seeds}}_{b}$ always dominates over $\mathrm{I}^{\textrm{resampling}}_{b}$, with the latter being between 0.2 and 0.5 times of the former on average. Therefore, training randomness contributes more to epistemic uncertainty than variability in data sampling. From a model development perspective, this indicates that if practitioners want to reduce their prediction uncertainty it may be more fruitful to train a bigger deep ensemble to target the uncertainty due to stochastic training, than to collect more data to target the uncertainty due to variability in data sampling.
Fig. 3 visualizes scatter plots of $\mathrm{I}^{\textrm{seeds}}_{b}$ versus $\mathrm{I}^{\textrm{deep ensemble}}_{b}$ along with the line of best fit (slope = $m_{\textrm{seeds}}^{\textrm{deep ensemble}}$) with numerical values of the slopes and correlation in Table 1. $\mathrm{I}^{\textrm{seeds}}_{b}$ and $\mathrm{I}^{\textrm{deep ensemble}}$ have similar magnitudes (around 90% of each other) and correlate well ($\rho\approx 0.9$). Therefore, deep ensembles do indeed capture the uncertainty due to training randomness. Together, these observations allow us to conclude that deep ensembles succeed by capturing the majority portion (training randomness) of an asymptotically correct estimator ($\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})$) of the Bayesian epistemic uncertainty.